# CALE : Concept-Aligned Embeddings for Both Within-Lemma and Inter-Lemma Sense Differentiation

### v1

@@ 6 Spatial Organization
In Figure 1 we plot the distributions of distance in pairs according to their categories: ‘SL’ and ‘DL’ refer to “same lemma” and “different lemmas” respectively, ‘SC’ and ‘DC’ to “same concept” and “different concepts” respectively. We first observe that XLM-R suffers from major anisotropy (i.e. all cosine distances are low even in unrelated pairs). The fine-tuning of CALE and XL-LEXEME push occurrences in DC/DL pairs away from one another, a consequence of the margin parameter. Another observation is that CALE’s fine-tuning successfully shifts the focus from lemma-centric to concept-centric: in XLM-R, DC/SL pairs (cases of polysemy) are generally closer than SC/DL pairs (cases of synonymy); this tendency is reversed in CALE (XLM-R) (average distance of .130.130 for DC/SL vs .157.157 for SC/DL before fine-tuning; .343.343 vs .269.269 after fine-tuning).
@@
WordNet’s synsets are linked to one another through relations of hyper-/hyponymy, in a tree-like structure. A short path between two synsets indicate that the two concepts are similar. We compute Spearman correlations between model’s cosine similarities and Wu-Palmer similarities (Wu and Palmer, 1994) between synsets for each occurrence pair [^5] We also tried Lin similarity, with similar results.. CALE models achieve the highest correlations (above 0.5), indicating a strong alignment between embedding similarity and semantic similarity. All correlations and pairwise differences are found statistically significant.
@@
We evaluate whether occurrences of the same concept form tight, well-separated clusters using the Silhouette score, with WordNet synsets as labels. CALE models yield the highest scores, suggesting better conceptual clustering. However, all scores remain below 0.1, highlighting substantial overlap between clusters even for the best models, which is likely due to the fine granularity of WordNet synsets.
@@ 7 Conclusion
Future work could explore specifically-curated datasets for Concept Differentiation, with human annotators and a different annotation scheme than WordNet’s synsets. We also look forward applications of CALE to Multiword Expressions (MWE), and potential applications to Frame Induction, Concept Induction or other semantic tasks.
@@ Acknowledgments
This research was funded by Inria Exploratory Action COMANCHE. We would like to thank Pascal Denis and Mikaela Keller for their valuable feedback on the early draft of this paper.
@@ Appendix D APD and PRT results on DWUGs
In addition to APD, we also tried another measure from Kutuzov and Giulianelli (2020), the Prototype Distance (PRT), defined as follows:

### v2

@@ 6 Spatial Organization
In Figure 1 we plot the distributions of distance in pairs according to their categories: ‘SL’ and ‘DL’ refer to “same lemma” and “different lemmas” respectively, ‘SC’ and ‘DC’ to “same concept” and “different concepts” respectively. We first observe that XLM-R suffers from major anisotropy (i.e. all cosine distances are low even in unrelated pairs of the DC/DL category). The fine-tuning of CALE and XL-LEXEME pushes occurrences in DC/DL pairs away from one another, a consequence of the margin parameter. Another observation is that CALE’s fine-tuning successfully shifts the focus from lemma-centric to concept-centric: in XLM-R, DC/SL pairs (cases of polysemy) are generally closer than SC/DL pairs (cases of synonymy); this tendency is reversed in CALE (XLM-R) (average distance of .130.130 for DC/SL vs .157.157 for SC/DL before fine-tuning; .343.343 vs .269.269 after fine-tuning). Yet, we remark that the cosine distances for several DC/SL pairs (examples of the polysemy of a single lemma) remain low: distances in the DC/SL category are not distributed in a tight mass, but rather a large band across the cosine distance range. We hypothesize that the reason for this low cosine distance value is that this corresponds to pairs of similar senses. Senses of a given lemma can be very similar to one another, and WordNet is known to be very fine-grained in its sense distinctions. Indeed, after examining this phenomenon more closely, we have found a (significant) correlation coefficient of 0.2 between Wu-Palmer similarity in WordNet and cosine distances between CALE (XL-LEX) embeddings. Thus, surprisingly low cosine distances in DC/SL pairs may be explained by similarity between word senses in the reference resource directly (at least partially, given that the correlation value is significant but not very high).
@@
WordNet’s synsets are linked to one another through relations of hyper-/hyponymy, in a tree-like structure. A short path between two synsets indicate that the two concepts are similar. We compute Spearman correlations between model’s cosine similarities and Wu-Palmer similarities (wu-palmer-1994-verb) between synsets for each occurrence pair [^5] We also tried Lin similarity, with similar results.. CALE models achieve the highest correlations (above 0.5), indicating a strong alignment between embedding similarity and semantic similarity. All correlations and pairwise differences are found statistically significant.
@@
We evaluate whether occurrences of the same concept form tight, well-separated clusters using the Silhouette score (Rousseeuw, 1987), with WordNet synsets as labels. Silhouette score measures how easily occurrences would be assigned to their cluster by comparing the distance to the occurrence’s cluster to the distance to the closest other cluster. A score close to 1 indicates that cluster are well-formed and well-separated. Scores close to 0 indicate that there exist overlaps between clusters, and scores close to -1 indicate that occurrences are almost systematically assigned to the wrong cluster. CALE models yield the highest scores, suggesting better conceptual clustering. However, all scores remain below 0.1, highlighting substantial overlap between clusters even for the best models, which is likely due to the fine granularity of WordNet synsets.
@@ 7 Conclusion
Future work could explore specifically-curated datasets for Concept Differentiation, with human annotators and a different annotation scheme than WordNet’s synsets. We also look forward to applications of CALE to Multiword Expressions (MWE), and potential applications to Frame Induction, Concept Induction or other semantic tasks.
@@ Acknowledgments
We gratefully thank the anonymous reviewers for their insightful comments. This research was funded by Inria Exploratory Action COMANCHE. We would like to thank Pascal Denis and Mikaela Keller for their valuable feedback on the early draft of this paper.
@@ Appendix D APD and PRT results on DWUGs
In addition to APD, we also tried another measure from kutuzov-giulianelli-2020-uio, the Prototype Distance (PRT), defined as follows: